media%2F634%2F634c3636-ccbb-4a1e-978b-01
Answer #1

At equilibrium:

Ag2CrO4 <----> 2 Ag+ + CrO42-

   2s s

Ksp = [Ag+]^2[CrO42-]

8*10^-12=(2s)^2*(s)

8*10^-12= 4(s)^3

s = 1.26*10^-4 M

Molar mass of Ag2CrO4,

MM = 2*MM(Ag) + 1*MM(Cr) + 4*MM(O)

= 2*107.9 + 1*52.0 + 4*16.0

= 331.8 g/mol

Molar mass of Ag2CrO4= 331.8 g/mol

s = 1.26*10^-4 mol/L

To covert it to g/L, multiply it by molar mass

s = 1.26*10^-4 mol/L * 331.8 g/mol

s = 4.18*10^-2 g/L

Now use:

volume = mass / solubility

= 3.5 g / 4.18*10^-2 g/L

= 83.7 L

Answer: 83.7 L